katherine:

If g(x) = x2 + 2, find g(3). 9 8 11 6

katherine:

@Felicity96

katherine:

is it 11?

Felicity96:

It looks like it is

katherine:

Thank you

katherine:

Can you help me on one more?

Felicity96:

Yea sure

katherine:

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(150)? This is the average number of days the house stayed on the market before being sold for $150,000. The house sold on the market for $150,000 and stayed on the market for an average of 150 days before being sold. The house sold for $150,000. The house stayed on the market for an average of 150 days before being sold.

katherine:

is it A?

Felicity96:

It sounds like its D.

katherine:

f(150) = f(p). This means that p=150. If p is the price in 1,000s, then f(150) = the number of days a house stays on the market before being sold for price 150,000$.

Felicity96:

Oh i guess i miss read the question. Im sorry

ThisGirlPretty:

@Arcadiouse

Arcadiouse:

Ok, what are we trying to solve here?

Felicity96:

Well see if im right on the first one then see who's right on the second one

Arcadiouse:

Yeah, I got 11 too.

Felicity96:

Me too

katherine:

what about the 2nd one?

katherine:

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(150)? This is the average number of days the house stayed on the market before being sold for $150,000. The house sold on the market for $150,000 and stayed on the market for an average of 150 days before being sold. The house sold for $150,000. The house stayed on the market for an average of 150 days before being sold.

katherine:

I think it's A

Arcadiouse:

Agreed

Arcadiouse:

Want to know the reason why?

Felicity96:

It is A?? I want to know the reason

Arcadiouse:

Yeah its a. p= price. and F(P) is equal to average number of days. So 150 thousand times f

Arcadiouse:

Theres no value for F, so you go off of what it tells you F(P) gives and what P is.

katherine:

So, is A correct?

Arcadiouse:

I would say so, yes.

Felicity96:

Oh ok

katherine:

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie. c(x) = 1.50 + 1.25x c(x) = 1.50x + 1.25 c(x) = 2.75 c(x) = (1.50 + 1.25)x

Felicity96:

B

Arcadiouse:

What do you think?

Arcadiouse:

Felicity, B is wrong. We also try to let them figure out the answer.

Felicity96:

Oh ok i didn't know that

Arcadiouse:

Ok, so if 1.50 is the flat rate and 1.25 is the nightly rate.

Arcadiouse:

What would you attach x/days too?

katherine:

Is it A?

katherine:

@Arcadiouse

Arcadiouse:

Correct.

katherine:

100 :O

katherine:

omg, thank you all so much!!!

Arcadiouse:

No problem! If you have any more questions feel free to mention me